OpenStudy (anonymous):

A wheelchair ramp rises 1.5 feet for every 18 feet of horizontal distance it covers. Find the rate of change of the ramp.

OpenStudy (baldymcgee6):

what's another word for rate of change?

OpenStudy (anonymous):

I don't know..

OpenStudy (baldymcgee6):



OpenStudy (baldymcgee6):

another word for rate of change is slope. slope = rise/run

OpenStudy (baldymcgee6):

\[\frac{ rise }{ run } = \frac{ 1.5 }{ 18 }\]

OpenStudy (anonymous):

Is the answer 16.5? ._.

OpenStudy (anonymous):

Nevermind. &Thanks.

OpenStudy (anonymous):

I think it's 0.83. Right? &Sure

OpenStudy (baldymcgee6):

close, its 0.083

OpenStudy (anonymous):

1/12